The obstacle problem for non local operators

Luis Caffarelli University of Texas at Austin, USA

We will dsiscuss the problem of inimizing the energy of a fractional derivative ( Hs) among those functions that must stay above a given obstacle: the motivation of the problem, regularity of the solution and the geometry of the free boundary.